OpenStudy (anonymous):

What is the area and perimeter of a square with the length of 1/3km and width of 1/8km?

OpenStudy (lgbasallote):

square?

OpenStudy (anonymous):

Uh..I thought squares had all the same sides....

OpenStudy (anonymous):

Area = L * W and Perimeter = 2(L) + 2(W) OR L + L + W + W <--4 sides total

OpenStudy (anonymous):

Area = 1/3km * 1/8km and Perimeter = (2)(1/3) + (2)(1/8)

OpenStudy (anonymous):

You know what I meant ^-^